### v3

Abstract
The advent of high-intensity radioactive ion beams has opened new avenues for nuclear structure research. By studying exotic ions, phenomena such as shell evolution, halos, and the limits of stability have been studied. In particular, Quasi-Free Scattering (QFS) experiments on hydrogen targets have proven to be a valuable tool to investigate the structure of exotic ions. Recently, a series of QFS experiments were performed at the Radioactive Isotope Beam Factory (RIBF) of the RIKEN Nishina Center employing the MINOS liquid hydrogen system. A fundamental part of the success of these experiments was the use of dedicated devices to measure the $\gamma$-rays, charged particles, and neutrons emitted in the reactions. The experience gained during the past campaigns, as well as the upcoming upgrade of the RIBF facility call for improvements on the existing devices, as well as for the development of new detection systems. Here we review the main detection devices used at the RIBF for QFS experiments, and give an overview of the ongoing and upcoming developments.
[…]
xxxx, xxx
1 Introduction
Recent advances in the production of Radioactive Ion Beams (RIBs) have transformed modern nuclear physics by allowing us to study the structure of exotic nuclei. This has provided us essential tests for models of the nuclear force and of nuclear reactions under extreme conditions Nowacki_PPNP_120_2021 . Furthermore, precision measurements of nuclear masses, half-lives, and reaction cross sections have improved astrophysical models of supernovae and neutron star mergers Aumann_PPNP_112_2020 . At the Radioactive Isotope Beam Factory (RIBF) of the RIKEN Nishina Center Yano_NIMB2007 , the unparalleled high primary beam intensities Accelerator2021_RAPR ; Accelerator2022_RAPR have been exploited to study phenomena such as the limits of nuclear existence Kondo_Nature_2023 ; Ahn_PhysRevLett.129.212502_2022 , shell evolution Steppenbeck_Nature_2013 ; Taniuchi_Nature_2019 , and multi-nucleon resonances Kisamori_PhysRevLett_116_2016 ; Duer_Nature_2022 .
Figure 1 shows the layout of the RIBF. Stable beams are accelerated up to 345 MeV/u using four cyclotrons. The last two, the Intermediate-stage Ring Cyclotron (IRC), and the Superconducting Ring Cyclotron (SRC), can be seen in Fig. 1. After exiting the SRC, beams are delivered to the F0 focal plane of the BigRIPS fragment separator Kubo_PTEP_2012 where RIBs are produced via in-flight fragmentation or fission on a thick target. Reaction products are then separated and identified event-by-event based on their mass-to-charge ratio, $A/Q$, and atomic number, $Z$ Fukuda_NIMA2013 .
Figure 1: Schematic of the RIBF facility. After the SRC, stable beam are delivered to the BigRIPS fragment separator where RIBs are produced, separated and identified. The different spectrometers available at the RIBF are shown, and different devices used for quasi-free scattering experiments are listed in red, at the focal plane where they are placed.
The resulting secondary beams, with energies from a few tens to several hundreds MeV/u, are delivered to one of the spectrometers of the RIBF: ZeroDegree Kubo_PTEP_2012 , SAMURAI Kobayashi_NIM_317_2013 , SHARAQ Michimasa_NIMA_317_2013 , OEDO Michimasa_PTEP_2019 , and the Rare-RI Ring YAMAGUCHI2013 , where different types of measurement can be performed. In particular, Quasi-Free Scattering (QFS) experiments have been mainly performed at the ZeroDegree and SAMURAI spectrometers. In these measurements RIBs with energies around 200 MeV/u impinge on a secondary target, leading to the knockout of one or few nucleons and approximating a free nucleon-nucleon collision. With this technique bound and unbound excited states, single-particle properties, and correlations within nuclei can be probed.
In recent years, a series of experiments involving QFS were performed at the RIBF employing the MINOS liquid hydrogen target Obertelli_EPJA_2014 ; santamaria2018tracking . Thanks to the time projection chamber (TPC) surrounding the target, it was possible to deduce the interaction vertex with a resolution of about 5 mm (FWHM) which, in turn, allowed for the use of a thicker target with higher luminosity. Different experiments with target lengths varying between 10 and 15 cm were performed at F8 and F13, and different detection systems were employed to measure the reaction residues.
At F8, the main focus was in-beam $\gamma$-ray spectroscopy following knockout reactions. In this case the DALI2+ array takeuchi2014dali2 ; Murray_RAPR_2018 was employed. This array, consisting of 226 NaI(Tl) detectors, provides high $\gamma$-ray detection efficiency which enables the study of bound excited states, as well as exclusive cross sections. At F13, MINOS was followed by the SAMURAI spectrometer to perform invariant mass measurements. In this case, in addition to the standard detectors of SAMURAI, which measure the $A/Q$ and $Z$ of the outgoing fragments, the neutron detectors NeuLAND Boretzky_NIM_2021 and NEBULA Kondo_NIM_2020 were employed. Furthermore, the DALI2+ array was also used in some experiments at F13 to measure excited states.
The combination of MINOS with these different detection systems was successfully used to measure phenomena such as magicity around 78Ni Taniuchi_PTEP_2025 , shell evolution around the Ca isotopes Lui_PTEP_2025 and the $N=40$ Island of Inversion (IoI) Cortes_PTEP_2025 , the limits of the $N=20$ IoI Kahlbow_PTEP_2025 , and the deformation of isotopes between Zn and Zr Chen_PTEP_2025 . These measurements also highlighted possible improvements to the detections systems and provided new ideas to further optimize the measurements. The continuous developments on the instrumentation is, therefore, of fundamental importance for studies involving QFS. This is particular the case at the RIBF, where an upgrade plan of the facility has been outlined for the coming decade ribf_upgrade .
In this article we review the main detection systems used for QFS experiments at the RIBF and give an overview of the on-going and future developments. Section 2 focuses on charged particle detectors for missing mass measurements, in particular STRASSE and TOGAXI. Section 3 gives an overview on the developments for in-beam $\gamma$-ray spectroscopy such as the HYPATIA array. Section 4 focuses on neutron detection arrays, in particular HIME and NEBULA+. In Sec. 5 a summary and outlook are presented.
2 Missing-mass spectroscopy in inverse kinematics
The nucleons in the atomic nucleus can be considered to be a system of independent particles within the Fermi gas model. The Fermi wave number, which defines the momentum of the highest occupied state in nuclear matter, is related to the typical nuclear saturation density of $\rho_{0}\approx 0.17~\mathrm{fm}^{-3}$. This value translates to $k_{F}=[3\pi^{2}/(2\rho_{0})]^{1/3}=1.36~\textrm{fm}^{-1}$, or approximately 270 MeV/$c$ in the momentum space. When the momenta of the particles significantly exceed Fermi momentum, the particles enter a quasi-free condition, no longer bound by the nuclear force. Therefore, proton-induced knockout reactions in inverse kinematics, expressed as $(p,pX)$, at beam energies above 200 MeV/u can be considered QFS reactions, where the proton scatters off a quasi-free cluster or nucleon inside the target nucleus.
((a))
((b))
Figure 2: (a) Dependence of recoil proton momentum on the scattering angle for different knockout reactions at 200 MeV/u. (b) Dependence of the QFS differential cross section on the momentum transfer ermisch2005systematic ; hasell1986elastic ; moss1980proton .
In $(p,pX)$ reactions the proton, initially at rest in the laboratory frame, gains momentum due to the reaction with the incoming heavy ion. Under the QFS approximation, it is assumed that the knocked-out particle acquires the same momentum transfer relative to the residual nucleus, as if the reaction was a direct, two-body collision, largely unaffected by the rest of the nucleus. Figure  2a) illustrates the dependence of the momentum transferred to the recoil proton as a function of the scattering angle. As the transferred momentum increases, the impact parameter between protons and intra-nuclear particles before the knockout becomes smaller. Consequently, the quasi-free scattering cross-sections tend to decrease, as shown in Fig. 2b).
Missing-mass spectroscopy following QFS reactions in inverse kinematics has become a key technique for investigating exotic nuclear structures, particularly in weakly bound and unbound nuclei, clustered systems, or exotic decays where direct detection is impractical. In this technique, undetected particles are identified by applying conservation of energy and momentum. Instead of detecting all final-state particles directly, the momenta and energies of the incoming beam and some of the outgoing particles are measured. The mass of the unobserved particle(s) is then deduced as the “missing” component needed to balance the reaction kinematics.
In particular for $(p,pX)$ reactions, the missing mass, $M_{\text{miss}}$, is deduced as
$M_{\text{miss}}=\sqrt{\left\{E_{b}+m_{p}-(E_{p}+E_{X})\right\}^{2}-\left\{\vec{P_{b}}-(\vec{P_{p}}+\vec{P_{X}})\right\}^{2}},$ (1)
where $E_{i}$, and $\vec{P_{i}}$ represent the energy and the three-momentum vector of the beam ($i=b$), the recoil proton ($i=p$) and the knocked-out particle ($i=X$), respectively, and $m_{p}$ is the proton mass.
((a))
((b))
Figure 3: Relation between emission angle and energy of (a) the scattered particles X, and (b) the recoil protons in $(p,pX)$ reactions at 200 MeV/u. The reaction Q-values are assumed to be zero for these calculations.
Figure 3 shows the relation between the energy, and emission angle, $\theta$, of a) the knocked-out particles (such as protons, deuterons, 3He, tritons, and $\alpha$ particles), and b) the recoil protons, for $(p,pX)$ reactions at incident energy of 200 MeV/u. For simplicity, the internal momentum of the knocked-out particle in the initial state is not taken into account. It can be seen that the mass of the knocked-out particles significantly influences $\theta_{X}$, particularly the maximum angles in which detection is possible: While knocked out protons can be detected at a larger angular range, heavier ejectiles are only emitted at forward angles, smaller than 30. In contrast, as can be seen on Fig. 3b), the recoil protons are scattered up to 90, regardless of the knocked-out particle.
When designing instrumentation for QFS reactions, it is essential to consider the angular distributions and energies of the scattered particles, as well as the cross sections required to achieve realistic yields. Detectors with different characteristics are therefore necessary to address different physics questions that can be studied with QFS reactions. The next subsections describe two major ongoing developments: STRASSE, optimized for nucleon knockout, and TOGAXSI, designed to study cluster knockout reactions. Each system is introduced with its specific motivation and improvement over existing approaches.
2.1 STRASSE: A telescope array for $(p,2p)$ and $(p,3p)$ reactions
STRASSE (Silicon Tracker for RAdioactive nuclei Studies at SAMURAI Experiments) strasse_epja is a detection system optimized for $(p,2p)$ and $(p,3p)$ reactions at 200–250 MeV/nucleon. Similarly to MINOS, a thick LH2 target with a length up to 150 mm is used to induce the QFS reactions. The recoil protons are detected, not by a TPC, but by a highly segmented Si tracker array placed in vacuum to minimize the angular straggling of protons. With this setup, proton detection efficiencies of 86% and 49% for one and two protons, respectively, can be achieved, comparable to those obtained with MINOS. Thanks to the high segmentation of the Si tracker, a vertex resolution of 0.5 mm (FWHM) and an angular resolution of $\sim 12$ mrad (FWHM), can be obtained.
STRASSE will be coupled with the CATANA (Cesium iodide Array for $\gamma$-ray Transitions in Atomic Nuclei at high isospin Asymmetry) array catana , composed of 140 CsI(Na) scintillators arranged into 7 rings (L2-L8), and covering polar angles from 17 to $7$. CATANA crystals can stop protons up to 250 MeV. In particular, a dual-gain readout has been developed for CATANA to enable simultaneous $\gamma$-ray and proton measurement. Note that the 20 crystals in L2 will be used to detect only protons. With the STRASSE + CATANA setup, missing-mass spectra, which could not be achieved with the MINOS system, are expected to be available with a resolution better than 4 MeV (FWHM). This offers an excellent opportunity to study highly unbound states (such as 4n or 6n states) or isomeric bound states (such as the $0_{2}^{+}$ states of doubly magic nuclei) at the RIBF.
[…]
Figure 4: Schematic view of the overall conceptual design of the STRASSE detection system.
[…]
2.2 TOGAXI: Elucidating cluster formation with $(p,pX)$ telescope
Clustering is a fundamental feature of nuclear structure, where nucleons tend to group into substructures such as $\alpha$ particles within a nucleus. By measuring cluster knockout reactions, it is possible to investigate the existence of clusters on the surface and inside of atomic nuclei. A notable example of $\alpha$-cluster formation in heavy ions was recently reported for the stable Sn isotopes Tanaka_Science_2021 . Analyzing the emitted particles after the $(p,p\alpha)$ reaction using a magnetic spectrometer, the cluster separation energy, a characteristic physical quantity of cluster knockout reaction, was deduced. Based on its agreement with theoretical models, a close connection between the $\alpha$-cluster formation and the neutron skin was suggested.
Similar types of studies have also been performed for exotic isotopes by using liquid or solid hydrogen targets and inverse kinematics. For example, in a recent measurement at the SAMURAI spectrometer, the measured triple differential cross section as a function of the proton energy revealed the existence of $\alpha$-clusters in 10Be, in agreement with a $\alpha-2$$n$ configuration matsuda2011large . Figure 5 shows the experimental setup used in that study. A 2-mm-thick solid hydrogen target (SHT) matsuda2011large was employed to induce $(p,p\alpha)$ reactions. The 6He residual nuclei were analyzed using the forward-located SAMURAI spectrometer. Because the emitted protons and alpha particles differ significantly in both emission angle and energy, separate detectors were placed to match each particle’s emission characteristics. Protons emitted at 60 degrees relative to the beam axis were measured with a recoil proton spectrometer (RPS-L(R)), composed of a drift chamber and a NaI (Tl) scintillator chebotaryov2018proton ; matsuda2013elastic , and $\alpha$ particles emitted between 4 and 12 degrees, were detected using a telescope consisting of a silicon detector and a CsI (Tl) scintillator (Tele-L(R)) verde2013farcos . By placing the detectors at a relatively large distance from the target, which minimizes the relative opening angle, an $\alpha$ separation energy resolution of 2.5 MeV (FHWM) was achieved. In addition, the time of flight of the emitted protons was measured.
[…]
In the cases described above, missing mass energy resolutions as low as 2.4 MeV (FWHM) has been achieved; however, the limited solid angle coverage remains a point for improvement. For next-generation detectors, it will be essential to inherit the successful concepts demonstrated by MINOS – namely, the use of a thick target and a detection system with large solid-angle coverage – to compensate for the low intensity of RI beams and to achieve high detection efficiency. Ideally, the system should be capable of measuring multiple types of cluster knockout reactions over a wide solid angle within a single configuration. Incorporating a total energy calorimeter for the emitted particles can further ensure accurate missing mass spectroscopy. The detector array introduced in the following paragraph is designed to simultaneously detect these charged particles, enabling a missing mass energy resolution of approximately 3 MeV (FWHM).
Figure 7 shows the TOGAXSI (TOtal energy measurement by GAgg and verteX measurement by SIlicon strips) telescope TANAKA20234 , being developed within the “ONOKORO” project uesaka2021comprehensive for dedicated measurements of cluster knockout reaction at the RIBF. Similarly to MINOS, it uses a thick LH2 target which provides high luminosity. For ONOKORO, the same type of liquid hydrogen target used for the STRASSE has been adopted. The target is surrounded by a Si tracker for vertex reconstruction and a GAGG:Ce calorimeter for energy measurement, which ensures accurate missing mass spectroscopy.
Figure 7: Graphical design of TOGAXSI array. Incoming beams interact in the LH2 target to produce cluster-knockout reactions. A Si tracker and a GAGG:Ce calorimeter are employed to measure the scattered clusters and the recoil protons and performe missing mass measurements.
[…]
As can be seen in Fig. 3a), under QFS conditions in inverse kinematics, the emitted clusters have high energies and are directed towards relatively forward angles, below 30, whereas recoil protons are emitted at larger angles. Consequently, the detector is designed to detect clusters at polar angles of $8^{\circ}$$30^{\circ}$. For the detection of recoil protons, detection angles between $35^{\circ}$$70^{\circ}$ are selected. As can be seen in Fig. 2, for these angles a momentum transfer between 2233 fm-1 is expected. This in turn leads to cross sections from $10^{-1}$ to $10^{-2}$ mb/Sr, which allows to keep a realistic yield for the experiments. In these angles the expected energy of the protons varies from 8080250250 MeV, and the cluster energy is between 7575180180 MeV/u.
The construction of the full array is underway. TOGAXSI is expected to accelerate comprehensive cluster research with missing mass spectroscopy based on the detection of charged particles. Table 1 summarizes the properties of MINOS, STRASSE and TOGAXI.
Table 1: Comparison of inverse-kinematics proton-induced reaction detector systems at RIBF.
Parameter MINOS Obertelli_EPJA_2014 STRASSE strasse_epja TOGAXSI TANAKA20234
Vertex resolution (FWHM) $\sim 5$ $m$ $\sim 0.5$ $m$ $\sim 0.5$ $m$
Angular resolution (FWHM) $\sim 38$ mrad $\sim 12$ mrad $\sim 10$ mrad
Missing-mass resolution (FWHM) $\leq 4$ MeV $\sim 3$ MeV
Angular coverage (cluster) $8^{\circ}$$30^{\circ}$
Angular coverage (recoil $p$) $14^{\circ}$$70^{\circ}$ $14^{\circ}$$70^{\circ}$ $35^{\circ}$$70^{\circ}$
3 Gamma-ray detection for gamma-ray spectroscopy
In-beam $\gamma$-ray spectroscopy is a powerful tool that allows the study of nuclear structure and excited states of atomic nuclei. The advent of RIBs has increased the potential of this technique as new mass regions can be explored and new phenomena can be observed. Indeed, high primary beam intensities have enabled the production of short-lived isotopes which in turn can impinge on solid or liquid targets to induce Coulomb excitation, inelastic scattering, knockout, and two-step fragmentation experiments in inverse kinematics Doornenbal_PTEP_2012 .
At the RIBF, secondary beams with incident energies of 100 to 250 MeV/u/u are used to populate excited states of nuclei far from stability. While such energies allow the use of thicker targets, and therefore increase the luminosity, they also present different experimental challenges. Of particular importance are the large Doppler shift of the detected $\gamma$-rays, and the background generated by either atomic or nuclear processes.
Observed $\gamma$-ray energies, $E_{\gamma}$, are Doppler shifted with respect to the emitted energies in the rest frame, $E_{\gamma 0}$, following the relation
$\frac{E_{\gamma 0}}{E_{\gamma}}=\frac{1-\beta\cos\theta_{\gamma}}{\sqrt{1-\beta^{2}}},$ (2)
where $\beta$ denotes the velocity of the beam and $\theta_{\gamma}$ is the angle between the particle direction and the $\gamma$-ray emission in the laboratory frame of reference. The energy resolution that can be achieved is therefore limited by the uncertainties in the velocity, $\Delta\beta$, the emission angle of the $\gamma$-ray, $\Delta\theta_{\gamma}$, and the intrinsic resolution of the detection material, $\Delta E_{\gamma}$. In general, the use of thicker secondary targets leads to a higher energy loss and therefore a larger contribution to $\Delta\beta$. The use of MINOS Obertelli_EPJA_2014 overcomes this problem by tracking the protons ejected from the target after knockout reactions and reconstructing the interaction position. In this way thicker targets can be used without affecting the achievable energy resolution as shown in Fig. 8. While the intrinsic resolution, $\Delta E_{\gamma}$, is a property of the detector material, the angular resolution, $\Delta\theta_{\gamma}$, can be improved either by reducing the size of the detectors or by implementing a position-sensitive readout with a sub-centimeter resolution for the first interaction of the $\gamma$-rays.
Figure 8: Total energy resolution (FWHM) for different $\gamma$-ray spectrometers. For a thick target, a large $\Delta\beta$ is obtained. This uncertainty is reduced by using MINOS. The use of smaller detectors with better intrinsic resolution, as the HYPATIA array, further improve the achievable energy resolution. The calculations were performed for $\beta=0.6$ and a 1 MeV $\gamma$-ray energy assuming a square-root dependence of the intrinsic energy resolution.
In addition to the energy resolution, the total efficiency of the detection array is an important factor. The use of high-efficiency materials becomes critical for the study of nuclei at the limit of existence where few events are expected, and for reconstructing complicated level schemes with multiple $\gamma$-ray emissions. In general, $\gamma$-ray detection arrays can be optimized either for an increased energy resolution or for a high detection efficiency.
A further parameter to considered is the time resolution. The high background generated in these type of experiments is usually a limiting factors, specially at low energies. It can be reduced by selecting events correlated to the impinging particle within a short time window. The successful use of this technique requires a time resolution typically better than 1 ns. Such resolution is only achievable for some scintillation materials.
The interplay of different factors makes it necessary to tackle the next generation $\gamma$-ray arrays from different perspectives, which depend on the physics cases pursued. In the following, we make a summary of future projects regarding $\gamma$-ray detection arrays at the RIBF.
3.1 Towards a new scintillation detector array: The HYPATIA project
Since 2001 the main device used at the RIBF to perform in-beam $\gamma$-ray spectroscopy has been the DALI2 array takeuchi2014dali2 , composed of 186 NaI(Tl) detectors. From 2017 the upgraded DALI2+ array Murray_RAPR_2018 , consisting of 226 NaI(Tl) detectors, covering a large solid angle and offering a high detection efficiency, has been employed to detect radiation emitted primarily after direct reactions with RIBs.
[…]
Figure 9: Conceptual design of the HYPATIA array composed of a barrel of CeBr3 detectors and a forward wall of HR-GAGG crystals.
Based on the experience acquired during the last two decades using DALI2 and DALI2+, the key characteristics desired for this new array are:
  • Higher detection efficiency which would allow to reach more exotic isotopes, as well as to perform detailed $\gamma-\gamma$ coincidences analysis to build level schemes.
  • Better energy resolution to identify previously unresolved transitions and exploit lifetime effects.
  • Improved time resolution to suppress the beam-related background.
  • A versatile design which allows to install of the array at different locations and with varying configurations optimized for each measurement.
With this requirement in mind, the HYPATIA (HYbrid Photon detector Array To Investigate Atomic nuclei) array has been designed hypatia_web . This array, depicted in Fig. 9, consists of 624 Cerium Bromide (CeBr3) detectors of $30\times 30\times 80$ mm2 forming a barrel around the target and 384 Cerium-doped Gadolinium Aluminum Gallium Garnet (HR-GAGG) detectors of $25\times 25\times 100$ mm2 placed as a forward wall. The CeBr3 detectors were chosen due to their improved energy and time resolution compared to NaI(Tl), their reasonable cost, and the lack of intrinsic radiation. On the other hand, the HR-GAGG crystals are non-hygroscopic and have a high effective atomic number, therefore offering the possibility to minimize the dead space and increase the efficiency and peak-to-total ratio at forward angles. Thanks to the cuboidal shape of all the detectors, it is possible to arrange them in different configurations, including surrounding the MINOS or STRASSE targets, as well as to place the array at the different spectrometers of the RIBF. Furthermore, the selected shapes and sizes of the detectors are compatible with the mechanical structure of the DALI2+ array, therefore it will be possible to gradually replace the detectors without interrupting the physics program.
The future availability of 238U primary beam at 345 MeV/nucleon with an intensity of 2000 pnA would allow the production of very exotic isotopes towards the drip lines. Inelastic scattering, Coulomb excitation, transfer reaction, and quasi-free scattering reactions can be employed to study these isotopes. Thanks to its superior efficiency, time resolution, and peak-to-total ratio, the HYPATIA array will be able to overcome some of the limitations of DALI2+, as well as of high-resolution Ge detectors which do not cover a solid angle of $4\pi$, such as the pandemonium effect. A key example of the possibilities of HYPATIA with the fast beam at the RIBF is the first spectroscopy of 60Ca.
Figure 10: a) Comparison between the performance of DALI2+ and HYPATIA for the measurement of 58Ca. b) Expected Doppler-corrected $\gamma$ spectrum of 60Ca for a 150-hours measurement with HYPATIA.
The Ca isotopic chain is particularly interesting for nuclear structure studies. While the proton shell closure at $Z=20$ seems to be maintained, new neutron shell closures at $N=32$ and $N=34$ have been evidenced by spectroscopy Huck_PhysRevC.31.2226 ; Steppenbeck_Nature_2013 , mass measurements Wienholtz_Nature2013 ; Michimasa_PhysRevLett.121.022506 , and knockout reaction cross sections Gade_PhysRevC.74.021302 ; Chen_PhysRevLett.123.142501 . The possibility of a further magic number at the $N=40$ harmonic-oscillator shell closure has been widely studied Gade_physics3040077 ; Brown_physics4020035 . Although measurements on 58Ca Chen_PLB_2023138025 and 62Ti Cortes_PLB_2020135071 , the closest neighbors of 60Ca, show no indication of a shell closure at $N=40$, only the first spectroscopy of 60Ca can provide definitive answers. This measurement will be possible with the HYPATIA array, as shown in Fig. 10b), where the simulated $\gamma$-ray spectrum obtained in a 150 h measurement is shown. As a reference, Fig. 10a) shows a comparison between the experimental spectrum obtained for 58Ca measured with DALI2+ Chen_PLB_2023138025 , and the expected performance of HYPATIA under the same conditions.
3.2 Future projects with Ge detectors
The $\gamma$-ray tracking technique of Ge detectors is the state-of-the-art method to achieve the best energy resolution for $\gamma$-rays. This technique has been implemented in the latest large-scale Ge arrays GRETINA PASCHALIS201344 and AGATA AKKOYUN201226 . During 2020 and 2021, the RIBF hosted the HiCARI array Wimmer2021 to conduct the first studies with tracking-type germanium detectors, which yielded successful outcomes. HiCARI was organized as a one-time campaign of experiments at the RIBF by gathering Ge detectors from other facilities. It is desirable to make continued efforts to push in-beam $\gamma$-ray spectroscopy based on tracking technology forward. For this reason, the Gamma-ray Tracking in R5 (GT-5) project was launched in 2023 at the RIKEN Nishina Center as a common and long-term platform for tracking-type Ge detectors at the RIBF to drive physics experiments and instrumentation for tracking technologies. The GT-5 has its own scientific and development programs related to in-beam $\gamma$-ray spectroscopy using RIBs. On a hardware basis, the GT-5 presently encompasses two in-house tracking detectors. One is the Compton camera telescope GREI motomura_IEEE_2007 , and the other is a position-sensitive strip germanium telescope SGT Suzuki_RAPR , both based on planar crystals with stripped segments. Furthermore, a GRETINA quad-type module was procured in 2023. In addition to the detectors, a maintenance facility for Ge detectors, electronics, and infrastructure is arranged. A unique aspect of the GT-5 is that it is not only involved in the scientific activities of the RIBF but also in the framework of the TRIP (Transformative Research Innovation Platform of RIKEN platforms) project, a RIKEN-wide initiative starting in 2023. One of the main goals of the TRIP use case is to establish an optical model for radioactive isotopes that are needed for reaction analyses and predictions but have yet to be studied based on systematic experimental data. To provide systematic data to optimize optical model potential parameters, the GT-5 is teamed with the elastic scattering measurement program MESA (Measurement of Elastic Scattering Anytime anywhere on any beam) and the reaction cross-section measurement program S3CAN (Symbiotic Systematic and Simultaneous Cross-section measurements for All over the Nuclear chart). As the first opportunities for data collection, experiments with stable beams and with RI beams have been conducted in 2024.
4 Neutron detection for invariant-mass spectroscopy
Invariant mass spectroscopy is a powerful tool for measuring the properties of nuclei under extreme isospin conditions. In particular, for neutron-rich isotopes, phenomena such as halo structures near the neutron drip line, neutron correlations, unbound states, and new cluster states are to be expected. The study of such properties allows us not only to probe the mechanisms underlying the nuclear force but also to better understand the nuclear equation of state and to investigate the properties of neutron stars. Such neutron-rich systems can be produced using different types of reactions with radioactive beams, and their structure can be determined by measuring the momenta of all the outgoing particles, including neutrons, following the decay of any intermediate resonance.
Kinematically complete measurements are performed at the SAMURAI spectrometer of the RIBF. For the detection of the fast neutrons evaporated after the reactions, the NEBULA (NEutron detection system for Breakup of Unstable nuclei with Large Acceptance) detector has been employed Kondo_NIM_2020 . NEBULA is made up of two identical wall structures, which consist of a layer of VETO scintillation detectors to eliminate hits from charged particles and two-layered neutron detectors with 30 each scintillation bars with readouts by PMTs at both ends. Each neutron detection bar is shaped as $12\times 12\times 180~\textrm{cm}^{3}$, featuring high neutron detection efficiency, large angular coverage, and capability of multiple neutron coincidence detection. A further improvement in the multiple neutron detection efficiency was achieved by combining NEBULA with NeuLAND (New Large-Area Neutron Detector), a new neutron detector array developed for the R3B setup under construction at FAIR Boretzky_NIM_2021 . With this combined setup, a diversity of phenomena has been measured, such as the first measurement of 28Kondo_Nature_2023 .
The upgrade plans of the RIBF will enable reaching of further neutron-rich isotopes, which are expected to have various weakly bound multi-neutron systems. To extend the sensitivity of multi-neutron reconstructions, a project to extend the detector volume for better detection efficiency, NEBULA Plus, has recently been completed through a collaborative effort between Japan and LPC Caen in France, as part of the French-funded EXPAND project expand . The NEBULA Plus array consisting of 90 new modules has been integrated with the existing array, effectively doubling the number of detection layers. This expansion increases the one-neutron detection efficiency by nearly a factor of 2 and improves the four-neutron detection efficiency by almost one order of magnitude. Initial in-beam tests were carried out at the RIBF in 2024 NebulaPlus_RAPR . Physics opportunities such as the investigation of the two-neutron decay of 21B, the search for 21B, and the measurement of a three-neutron decay of 20B, are expected to be accessible with the NEBULA Plus detector.
In contrast, a detector focused on more precise measurement by improving the achievable resolution in measurements of neutron momentum is also being constructed. The High-resolution Detector Array for Multi-neutron Events (HIME) detector was originally developed by the group of The Tokyo Institute of Technology (now Institute of Science Tokyo), and further developed in collaboration with the Technical University of Darmstadt. The optimized energy resolution of HIME is achieved through the use of smaller scintillator bars with dimensions of 4 $\times$ 2 $\times$ 100 cm3, which improves position resolution, and new electronics for a higher time precision. It consists of layers of scintillator bars arranged alternating in $x$ and $y$ directions Knoesel_thesis . This arrangement will not only improve its resolution, but will also allow discriminating the hits due to recoiling protons, allowing the selection of valid events and reducing cross-talk Hime_report . Currently, a prototype of HIME has been built, and more layers are under construction at TU Darmstadt. The improved energy resolution will enable precise measurements, such as the neutron-neutron scattering length, which characterizes the two-neutron interactions at low energies. Besides that, it is planned to be used for the detection of the recoil neutrons after the $(p,pn)$ reactions in combination with the proton trackers as described in the previous sections.
5 Summary
Thanks to advances on the production of RIBs, isotopes that were previously inaccessible have been studied, and many interesting phenomena have been discovered. An integral part of research on the structure of exotic isotopes is the use of appropriate devices to detect the $\gamma$-rays, charged particles, and neutrons produced in the reactions.
A successful campaign of QFS experiments was performed at the RIBF using the MINOS LH2 target. The detection devices used for these experiments provided a plethora of results, and highlighted the limitations and possible improvements of the different detection systems. This experience, together with the future upgrade of the RIBF have promoted the development of new and improved detectors.
Based on the good performance of MINOS, the STRASSE Si tracker is under development. In addition to an improved resolution on the vertex reconstruction for knockout reactions, this system will provide the possibility to perform missing-mass measurement by the coupling with CATANA PLUS. A similar development, TOGAXI, will enable dedicated measurements of cluster knockout reaction. Thanks to the use of a calorimeter, missing mass-measurements will also be possible. Invariant mass measurements will benefit from the development of the neutron detectors NEBULA Plus and HIME, optimized for an increased efficiency and resolution, respectively. Furthermore, the study of bound excited states will profit from the developments on $\gamma$-ray detections arrays, such as HYPATIA, which will provide a better energy and time resolutions, as well as an improved peak-to-total ratio. Applications requiring optimized energy resolution can be addressed within the framework of the GT5 project.
In the future, the combination of QFS experiments with these new detection devices will help us to broaden our understanding of the structure of exotic isotopes. These detectors will also provide benchmarks for future developments in combination with the upgraded intesities of the RIBF.
Acknowledgment
The authors acknowledge the contributions of Pieter Doornenbal, Marco Knösel, Alexandre Obertelli, Nigel Orr, Valerii Panin, Marina Petri, Daisuke Suzuki, Yasuhiro Togano, Ryotaro Tsuji, Tomohiro Uesaka, and Rin Yokoyama. J.T. acknowledges the ONOKORO collaboration and the support from the Grants-in-Aid for Scientific Research of the Japan Society for the Promotion of Science (JSPS) KAKENHI, Grant No. JP21H04975. M.L.C acknowledges the support of the RIKEN Special Postdoctoral Researcher Program. H.N.L thanks the STRASSE-CATANA collaboration, and acknowledges the support from the National Key R&D Program of China (Grant No. 2023YFA1606403, No. 2024YFE0102800), the National Natural Science Foundation of China (Grant No. 12375111), and the Fundamental Research Funds for the Central Universities of China. R.T acknowledges the support from the UK STFC and the Royal Society.